HIGH RESOLUTION INVESTIGATION OF THE ETHANE SPECTRUM AT 7 MICRON (1430 CM^{−1})

C. DI LAURO, F. LATTANZI, Dipartimento di Chimica Farmaceutica e Tossicologia, Università di Napoli Federico II , I-80131 Naples, Italy; K. SUNG, \underline{L. R. BROWN}, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109, USA; J. VANDER AUWERA, Service de Chimie Quantique et Photophysique, Université Libre de Bruxelles, CP 160/09, 50 avenue F.D. Roosevelt, B-1050 Brussels, Belgium; A.W. MANTZ, Dept. of Physics, Astronomy and Geophysics, Connecticut College, New London, CT 06320, USA; M. A. H. SMITH, Science Directorate, NASA Langley Research Center, Hampton, VA 23681, USA.

Building upon our previous contributions,^{a} we are re-investigating the ethane spectrum between 1330 and 1610 cm^{−1}. For this, spectral data were obtained at room and cold (130 K) temperatures with two Bruker Fourier transform spectrometers (at 0.002 cm^{−1} resolution in Brussels and at 0.003 cm^{−1} resolution in Pasadena). Over 3300 lines were assigned to ν_{6}, ν_{8}, ν_{4} +ν_{12} and 2ν_{4} +ν_{9} cold bands, and one hot band (ν_{4} +ν_{8} −ν_{4}). Note that ν_{6}, ν_{8}, ν_{9}, and ν_{12} are near 1379, 1472, 823, and 1195 cm^{−1}, respectively, and ν_{4} is the torsional mode near 289 cm^{−1}. Our new analysis includes an improved implementation of the theoretical Hamiltonian needed to interpret the very complex spectral structures caused by numerous interactions between these 5 vibrational modes. From this, an empirical line list of positions and estimated intensities is being generated for planetary applications.

Part of the research described in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, Connecticut College, and NASA Langley under contracts and grants with the National Aeronautics and Space Administration.

aF. Lattanzi, C. Di Lauro, M. Herman, J. Vander Auwera, J. Mol. Spectrosc. 216 (2002) 308-314; F. Lattanzi, C. Di Lauro, V.-M. Horneman, M. Herman, J. Vander Auwera, Mol. Phys. 105 (2007) 733-740; F. Lattanzi, C. Di Lauro, J. Vander Auwera, J.

Mol. Spectrosc. 248 (2008) 134-145.